OpenStudy (anonymous):

Write the equation of the circle, given the following conditions Center (7,-4) Radius 6 units a. (x-7)^2 + (y+4)^2 = 36 b. (x+7)^2 + (y+4)^2 = 36 c. (x-7)^2 + (y+4)^2 = 6 d. (x-7)^2 + (y-4)^2 = 36 How do I do this?

OpenStudy (anonymous):

(x-7)^2 + (y+4)^2 = 36

Directrix (directrix):

You need to know the standard form for the equation of a circle so that you can identify the center and radius. (x -( h)) ^2 + (y -( k))^2 = r^2 h = 7, k = -4, and r = 6 (x - 7) ^2 + (y - (-4) )^2 = 6^2 (x - 7) ^2 + (y + 4) )^2 = 36 --> Option A